# Vid2Robot: End-to-end Video-conditioned Policy Learning with Cross-Attention Transformers

### v1

@@ II-D Preprocessing
To handle the varying lengths of videos for efficient training, we randomly sample N=16N\!\!=\!\!16 frames always including first and last frames and sort them in increasing order of time. During training, we sample a robot state $S_{t}$ by sampling a random timestep first. We then select the preceding k1k-1 frames to create a robot state video comprising of a total of k=8k\!\!=\!\!8 frames before. In case there are less than k1k-1 frames before the current time-step, we repeat the first frame to create a fixed size robot state video. The pixel values in each frame are normalized between 0 to 1. Each frame is resized to $(224,224)$. Photometric distortions like cropping, brightness, contrast, hue and saturation are applied during training.
The action vector consists of values indicating the mode, gripper pose and closedness as well as base displacement and rotation. Each of the values have different ranges, which we first use to scale the values in between 0 and 1. We then discretize the values into 256 bins each. In total, we construct 11-dim action vector as target, each of which has value between [0, 255]. In this study, we train and evaluate in the scenarios where base remains stationary.
@@ II-E Training
Figure 4: Training Setup. We show all the losses Vid2Robot is trained with and how each loss is connected to its different modules. Along with (1) the main action prediction loss, we apply three auxiliary losses: (2) temporal video alignment loss, (3) a contrastive loss between the prompt and robot video performing the same task, and (4) a contrastive loss between a prompt/robot video with the language embedding. More details in §II-E.
Action Prediction Loss We train Vid2Robot end-to-end with behavior cloning. The idea is to learn video representations from raw pixels to recognize task verb and objects, as well as learn motor control to accomplish it. We use a classification loss on actions that have been tokenized into N=256N\!\!=\!\!256 bins. Given the robot trajectory for performing a task with current visual observations $x_{t}$, we have the corresponding expert action $a_{t}$. The action prediction loss is Cross Entropy between the predicted action and the expert action as:
$L_{CE}(a_{t},\hat{a}_{t})=\sum_{\tau}a_{t}\log\hat{a}_{t}$ (1)
This trains all the model parameters, as shown in Fig 3.
Auxiliary Losses. Although our dataset size is substantial, it is insufficient for training large Transformer based models. In order to prevent over-fitting by just predicting actions correctly on the training set, we add three auxiliary losses that encourage learning features that are helpful in understanding semantics in prompt videos.
Video Alignment Loss: We want to encourage temporal alignment between prompt videos and robot videos performing that show the same task. By aligning prompt videos and robot videos, we want the image encoder to learn to be invariant to different embodiments, lighting, backgrounds, view-angles and distractor objects while still encoding features relevant to predicting task progress. Our choice of loss is the temporal-cycle consistency loss introduced in [17]. This loss has been shown to encode task progress when trained on videos of different agents performing the same task [47]. This loss is applied on per-frame image embeddings of the prompt $V_{p}$ and robot $V_{r}$ videos during training. To apply the loss, we average pool the per-frame embeddings output in spatial dimensions from image encoder $\phi$ and apply a projector head of 2-layer MLP [10]. We call this as alignment pooling layer $\Phi$ on the per-frame image embeddings, as shown in Fig 4. For each video $V_{i}$, this results in a sequence of embeddings $E_{i}=\{\Phi(v_{i}^{1}),\Phi(v_{i}^{2})...,\Phi(v_{i}^{L_{i}})\}$, where $L_{i}$ is the length of the $i^{th}$ video.
We apply TCC loss on encoding $E_{p}$, and $E_{r}$ for prompt and robot video respectively. The intuitive idea of TCC loss is that we want to ensure the representation of every frame of $E_{p}$ should have a correspondence in $E_{r}$ and vice versa. This involves two steps: First, we compute soft neighbor of $t^{th}$ frame of $E_{p}$ ($E_{p}^{t}$ in short) in $E_{r}$ and call it $\widetilde{E_{pr}^{t}}$.
@@
Second, we find the corresponding frame for this newly computed soft-neighbour in $E_{p}$. This is called cycle-back in [17] and it involves similar soft-neighbour computation as in Equation 2 to obtain say $\widehat{E_{pr}^{t}}$, which ideally should be same as $t$, that is, $(\widehat{E_{pr}^{t}}-t)^{2}$ should be minimized. TCC loss minimizes such mean squared error between all frames for prompt and robot video encodings and vice-versa, that is,
@@
Prompt-Robot Video Contrastive Loss (VVCL): We apply contrastive loss between prompt tokens produced by robot or prompt video performing the same task. This loss encourages the prompt encodings to learn task semantics from video tokens only in a self-supervised manner. A thing to note here is that the initial pairing of prompt and robot video has been done using natural language. However, by using a constrastive loss only on video embeddings with a self-supervised loss, we hope to encode features not covered by the short natural language embedding itself. Examples of these features include similar motions like reaching for objects, and rotating the robot arm. We use a Attention Pooling layer to merge features from the $N$ prompt tokens to produce a single embedding for each video. We apply the SigLIP [48] loss between video-video pairs to encourage videos showing same task, involving similar motions and interacting objects, to be close to each other while being away from other videos in the batch. A batch contains the same number of robot videos and prompt videos, say $B$. We use the prompt encoder $\psi_{p}(\phi(\cdot))$ to obtain a batch of full robot video embeddings $Z_{robot}$ and prompt video embeddings $Z_{prompt}$, each of size $B\times d$. We multiply them, $Z_{robot}\cdot Z_{prompt}^{T}$ to obtain a $B\times B$ matrix. Adding a learnable temperature $\tau$ and bias $b$, we have our logit matrix as $\hat{Y}=(Z_{robot}\cdot Z_{prompt}^{T})*\tau+b$. We consider the videos of robot and prompt performing the same task as positives and assign them a label of 1 along the diagonal and -1 for off-diagonal pairs, that is, the label matrix $Y=2\mathrm{I}_{B}-1$. SigLIP loss is the negative loglikelihood, specifically, $\sigma^{\prime}(Z_{1},Z_{2})=-\sum\log\sigma(Y\cdot(Z_{1}\cdot Z_{2}^{T})*t+b)\text{, where }Y=2\mathrm{I}_{B}-1$. The video-video contrastive loss is then defined as:
@@
Video-text Contrastive Loss (VTCL): We add a contrastive loss between prompt tokens $Z_{prompt}$ and those produced by robot video $Z_{robot}$ and the embedding of the text instructions of the task $Z_{text}$. This encourages a part of the embedding space to be aware of object names and verbs present in the prompt and the robot videos. A version of this loss has been applied before by BC-Z [21] as auxiliary language regression loss. We use an Attention Pooling layer [46] with one latent query to merge features from the $N$ prompt tokens to produce a single embedding for each video. Within a batch, we retrieve $B$ pairs of video and text embeddings. Similar to Equation 4, we apply SigLIP [48] loss to get
$L_{VTCL}=\frac{\sigma^{\prime}(Z_{prompt},Z_{text})+\sigma^{\prime}(Z_{robot},Z_{text})}{2}$ (5)
This encourages every video to have similar embeddings to their textual description embeddings, while being different from the text embeddings corresponding to other videos in the batch.
@@ II-F Implementation
We trained the model (implemented in Jax) for 200K iterations. We use AdamW optimizer with an initial learning rate of 8e-5 using a cosine learning rate schedule with warmup steps 2,000 and final learning rate of 1e-6. For both the Prompt and State Resamplers, we use 2 Perceiver Resampler layers with 64 latents. Both state-prompt encoder and action decoder are 4 layer deep cross-attention transformers.
@@ III Experiments
Figure 5: Policy Rollouts. Each row shows a prompt video of a human doing a task on the left, and on the right we show the corresponding successful robot rollouts using Vid2Robot. Note how visually different the prompts are, while the policy rollouts are recorded with different lighting, background, as well as number and placement of the distractor objects.
We present results with real robot evaluations for our multi-task video-conditioned policy. One of the key questions that we tackle in this work is how well robots can imitate humans performing manipulation tasks. Because of differences in embodiments, humans perform manipulation tasks at a different speed and style. We study the effect of using robot as well as human videos as prompts.
Metrics. We refer to a rollout as a sequence of actions inferred from the policy and executed on the robot from an initial state observation and prompt video, until the policy terminates or a maximum number of steps are taken, whichever is lower. We define success for a rollout when the policy executes the task instruction shown in the prompt video. A successful rollout involves correct actions to be taken successively in the environment, without any assistance for resets or recovery. For each task instruction, we record many rollouts per policy. We take the average of success recorded across all the rollouts and call it the Success Rate for that task. Aggregated success rate across tasks is referred as Overall Success Rate.
A mistake made early on in a rollout can result in poor success rate, even if the model’s offline overall prediction accuracy is high. For example, if a policy makes an error while grasping a water bottle early on in the task and it slips to an unreachable location, the rollout will be marked as a failure, even if the policy had good action predictions for the later steps. To record partial progress for a rollout, we annotate whether the robot reached the correct location, grasped the correct object, released the object at the correct location, and terminated the task correctly. More details on partial success analysis in §III-A.
Evaluation Setup. We ask human raters to evaluate success for a policy’s rollout on a robot. We evaluate the policies by varying the robots, lighting conditions, chest of drawers, and objects. We ensure the policies being evaluated are shown similar initial object configurations during rollouts. The initial state is randomized after all policies have been evaluated for a given initial state. For all rollouts, we sample prompt videos that are not seen by the models during training. This ensures that the policies are evaluated for whether they can recognize the task from new prompt videos or not.
Baselines. We compare our model with BC-Z [21], a video conditioned policy using a ResNet-18 encoder. BC-Z [22] involves demonstration-observation pairs processed via a FiLM [36] conditioned ResNet encoder and fed into a ResNet based policy network to predict robot actions. For a fair comparison, we train the BC-Z model with the same training data used to train the Vid2Robot model. We run rollouts of BC-Z policy for a fixed maximum number of steps, as BC-Z doesn’t have a terminate action.
@@
Key Questions and Results We address the following questions in this work:
  1. 1.
    How well do video-conditioned policies perform when they are shown a task in an unseen video? (Fig 5, § III-A)
  2. 2.
    What is the gap in success rate due to prompt embodiment difference (robot v/s human)? (§ III-A)
  3. 3.
    Can we leverage the learned motion representations for out-of-distribution object interactions? (§ III-B)
@@ III-A Task-based success
We compare the our Vid2Robot model and baseline BC-Z with robot and human prompt videos in Table I. Both Vid2Robot and BC-Z were trained on a same data mixture containing robot-robot and human-robot paired data. Prompt videos cover a subset of the training tasks but the videos themselves are new for the models. In this evaluation, we investigate what each model’s ability is to infer the task specification from prompt video as well as the current observed state of the robot.
In order to test the capabilities of the model in different settings on real robot, we evaluate it across eight categories of manipulation tasks as shown in Table I. Specifically, we evaluate for nine tasks: ‘knock water bottle over’, ‘move rxbar chocolate near coke can’, ‘move green jalapeno chip bag near coke can’, ‘pick green rice chip bag’, ‘place coke can upright’, ‘pick coke can from bottom drawer and place on counter’, ‘open middle drawer’, ‘close middle drawer’, and ‘place apple into top drawer’.
We ask four evaluators to carry out two rollouts per task for a prompt video dataset and policy setting (a row in Table I), that implies, we have eight trials per task to evaluate a policy’s task success rate. We report overall success rate per row over nine tasks with eight trials per task, that is, $9\!\times\!8\!=\!72$ trials. In total, our evaluations in Table I required $72\!\times\!4\!=\!288$ real robot rollouts.
@@ III-A1 What is the gap in success rate due to embodiment difference in prompt videos?
We compare our model with BC-Z when prompted with robot and human videos. BC-Z serves as a strong baseline for our comparisons. The overall success rate of our model Vid2Robot outperforms BC-Z for Human prompt videos by 20%, and is comparable for Robot prompt videos. Note that there is an order of magnitude more training samples for robot trajectories than human videos in our training mixture. Hence, there isn’t a significant gap in performance for robot prompt videos. For human prompt videos, our model outperforms BC-Z in most tasks, showing that Vid2Robot captures the task semantics from prompt videos better than the baseline. Our model outperforms in tasks like picking from drawer and placing on the counter, and opening/closing drawer tasks by a large margin. The most challenging task is placing upright and knocking over. We analyze the failure reasons in §V Fig 9.
III-A2 How well do video-conditioned policies perform when they are shown a task in an unseen video?
In addition to marking a rollout as a success, we recorded partial success annotations per rollout. In Fig 6, we observe that our model reaches to the correct object 78% of the time, about 8% more as compared to baseline. The policies sometimes fail to reach the correct object and go towards a distractor instead. Next, grasping errors happen, particularly with small and deformable objects and in collision prone areas like drawer handle or counter’s edge. Here our model (65%) outperforms BC-Z (45%) by a large margin of 20%. A successful grasp often the most difficult part in a rollout, and the most crucial for success. After grasping, most tasks require releasing at a correct location. There is a slight drop in success rate in both models due to incorrect release during the rollouts. While BC-Z runs for a fixed number of steps, our policy Vid2Robot predicts when to terminate. We observe that the rate of release and terminate is almost identical, about 57% for our model, that implies, that after releasing at correct location, Vid2Robot mostly terminates successfully.

### v2

@@ II-D Preprocessing
To efficiently train videos of varying lengths, we randomly sample $N\!\!=\!\!16$ frames. We always include the first and last frames and sort them in increasing order of time. We sample a robot state $S_{t}$ during training by sampling a random timestep. We then select the preceding $k-1$ frames to create a robot state video comprising a total of $k\!\!=\!\!8$ frames before. If there are less than $k-1$ frames before the current time step, we repeat the first frame to create a fixed-size robot state video. We normalize the pixel values in each frame between 0 and 1 and resize each frame to $(224,224)$. During training, we apply photometric distortions like cropping, brightness, contrast, hue, and saturation.
The action at that time consists of the three components: Mode: ($m$) whether to terminate episode, move only arm, move only base, or both. arm: gripper position (x, y, z), orientation (rotation along xy, yz, zx), and the degree of closedness (c). Base: displacement (x, y) and rotation. Overall, the action $a_{t}=[m,g_{x},g_{y},g_{z},\theta_{xy},\theta_{yz},\theta_{zx},c,b_{x},b_{y},b_{\theta}]$ is an 11-dimensional vector. Each value has different ranges, which we first use to scale the values between 0 and 1. We further discretize the values into 256 bins each. In this study, we train and evaluate scenarios where the base remains stationary.
@@ II-E Training
Figure 4: Training Setup. We show all the losses used for training Vid2Robot, particularly how each loss connects to its different modules. Along with (1) the main action prediction loss, we apply three auxiliary losses: (2) temporal video alignment loss, (3) a contrastive loss between the prompt and robot video performing the same task, and (4) a contrastive loss between a prompt/robot video with the language embedding. More details are in Section II-E.
(1) Action Prediction Loss. We train Vid2Robot end-to-end with behavior cloning. We use a classification loss on actions discretized into $N\!\!=\!\!256$ bins. Given the robot trajectory for performing a task with current visual observations $x_{t}$, we have the corresponding expert action $a_{t}$. The action prediction loss is Cross Entropy between the predicted action and the expert action as: $L_{CE}(a_{t},\hat{a}_{t})=\sum_{\tau}a_{t}\log\hat{a}_{t}$. This loss trains all the model parameters, as shown in Fig 3.
Although our dataset size is substantial, it is insufficient for training large transformer-based models. To prevent over-fitting on the training set, we add three auxiliary losses to encourage learning features in prompt videos.
(2) Video Alignment Loss: We want to encourage temporal alignment between prompt videos and robot videos performing that show the same task. By aligning prompt videos with the robot videos, we want the image encoder to learn to be invariant to different embodiments, lighting, backgrounds, view angles, and distractor objects while still encoding features relevant to predicting task progress.
Our choice of loss is the temporal-cycle consistency loss introduced in [18]. This loss can encode the task progress when trained on videos of different agents performing the same task [51]. This loss is applied on per-frame image embeddings of the prompt $V_{p}$ and robot $V_{r}$ videos during training. To apply the loss, we average pool the per-frame embeddings output in spatial dimensions from image encoder $\phi$ and apply a projector head of 2-layer MLP [11]. We call this as alignment pooling layer $\Phi$ on the per-frame image embeddings, as shown in Fig 4. For each video $V_{i}$, this results in a sequence of embeddings $E_{i}=\{\Phi(v_{i}^{1}),\Phi(v_{i}^{2}),...\Phi(v_{i}^{L_{i}})\}$, where $L_{i}$ is the number of frames in the $i^{th}$ video. We apply TCC loss on encoding $E_{p}$ and $E_{r}$ for prompt and robot video, respectively. Intuitively, the TCC loss ensures that the representation of every frame of $E_{p}$ should correspond to $E_{r}$ and vice versa. Applying TCC in Vid2Robot involves two steps: First, we compute soft neighbor of $t^{th}$ frame of $E_{p}$ ($E_{p}^{t}$ in short) in $E_{r}$ and call it $\widetilde{E_{pr}^{t}}$.
@@
Second, we find the corresponding frame for this newly computed soft-neighbor in $E_{p}$. This is called cycle-back in [18] and it involves similar soft-neighbour computation as in Equation 1 to obtain say $\widehat{E_{pr}^{t}}$, which ideally should be same as $t$, that is, $(\widehat{E_{pr}^{t}}-t)^{2}$ should be minimized. TCC loss minimizes such mean squared error between all frames for prompt and robot video encodings, and vice-versa, that is,
@@
(3) Prompt-Robot Video Contrastive Loss (VVCL): We want to encourage the prompt encodings to learn task semantics from video tokens in a self-supervised manner. While we pair prompt and robot video using natural language, this does not effectively capture the visual similarity of low-level motions like reaching for objects and rotating the robot arm. For this, we apply contrastive loss between the latent features of the robot and the prompt videos. We use an Attention Pooling layer to merge features from the $N$ prompt tokens to produce a single embedding for each video. We apply the SigLIP [52] loss between video-video pairs to encourage videos showing the same task, involving similar motions and interacting objects, to be close to each other while being away from other videos in the batch. A batch contains the same number of robot and prompt videos, say $B$. We use the prompt encoder $\psi_{p}(\phi(\cdot))$ to obtain a batch of full robot video embeddings $Z_{robot}$ and prompt video embeddings $Z_{prompt}$, each of size $B\times d$. We multiply them, $Z_{robot}\cdot Z_{prompt}^{T}$ to obtain a $B\times B$ matrix. Adding a learnable temperature $\tau$ and bias $b$, we have our logit matrix as $\hat{Y}=(Z_{robot}\cdot Z_{prompt}^{T})*\tau+b$. We consider the videos of robot and prompt performing the same task as positives and assign them a label of 1 along the diagonal and -1 for off-diagonal pairs, that is, the label matrix $Y=2\mathrm{I}_{B}-1$. SigLIP loss is the negative loglikelihood $\sigma^{\prime}(Z_{1},Z_{2})=-\sum\log\sigma(Y\cdot(Z_{1}\cdot Z_{2}^{T})*t+b)$. The video-video contrastive loss is as follows:
@@
[…]
$L_{VTCL}=(\sigma^{\prime}(Z_{prompt},Z_{text})+\sigma^{\prime}(Z_{robot},Z_{text}))/2$ (4)
@@ II-F Implementation
We trained the model (implemented in Jax) for 200K iterations. We use AdamW optimizer with an initial learning rate of 8e-5 using a cosine learning rate schedule with warmup steps 2,000 and a final learning rate of 1e-6. We use 2 Perceiver Resampler layers with 64 latent tokens for both the Prompt and State Resamplers. Both state-prompt encoder and action decoder are 4-layer deep cross-attention transformers.
@@ III Experiments
Figure 5: Policy Rollouts. Each row shows a prompt video of a human doing a task on the left, and on the right, we show the corresponding successful robot rollouts using Vid2Robot. Note how visually different the prompts are, while the policy rollouts have different lighting and backgrounds, as well as the number and placement of the distractor objects.
@@
We present results with real robot evaluations for our multi-task video-conditioned policy. One of the fundamental questions we tackle in this work is how well robots can imitate humans performing manipulation tasks. Because of differences in embodiments, humans perform manipulation tasks at a different speed and style. We study the effect of using robots and human videos as prompts.
Metrics. A rollout as a sequence of actions inferred from the policy and executed on the robot from an initial state observation until the policy terminates or takes the maximum number of steps, whichever is lower.
We define success for a rollout when the policy executes the task instruction as shown in the prompt video. A successful rollout involves correct actions to be taken successively in the environment without any assistance for resets or recovery. We ask a human evaluator to observe whether: (1) whether the robot reached the correct location?, (2) grasped the correct object?, (3) released the object at the correct location?, and (4) terminated the task correctly?
If the answer to any question is “no”, the answer to subsequent questions is assumed to be “no”. If all questions are answered “yes”, only then the rollout is considered “successful”. For each task instruction, we record a few rollouts per policy with different distractor objects, background, and lighting conditions. We take the average success recorded across all the rollouts for a task and call it that task’s Success Rate. We also report aggregated success rate across tasks as Overall Success Rate. We also analyze the partial success across the four milestones in Section III-A.
We looked into automating success detection with concrete criteria or decision rules for evaluation but found that this automation can overlook the process of performing the task. Consider the task of “knocking the water bottle over”. Case 1: The robot successfully grasps the bottle, turns it, and places it on the table. Case 2: The robot fails to grasp, pinches the bottle away, and lands in a knocked-down orientation. While Case 1 is desired and expected behavior according to the training data, Case 2 is a failure as it is unintended. With rule-based verification of the final state, we would have deemed both Case 1 and 2 successful. With human evaluators, we focus on the entire process of achieving the task instead of the final state.
Setup. We evaluate the policies by varying the object placement, lighting conditions, background surfaces, and distractor objects. When evaluating a set of policies, we ensure comparable initial object configurations for rollouts. We randomize the initial state only after all policies’ rollouts have been performed. For all rollouts, we sample prompt videos not seen during training. In all the experiments, we use a mobile manipulator, the Google Robot. It has an ego-centric camera view, a single arm with seven degrees of freedom, and a two-fingered soft gripper. Refer to [8] for more details.
Baselines. We compare our model with BC-Z [22], a video-conditioned policy using a ResNet-18 encoder. BC-Z [23] processes demonstration-observation pairs via a FiLM [38] conditioned ResNet encoder and feds into a ResNet-based policy network to predict robot actions. We use the same data to train the BC-Z and Vid2Robot for a fair comparison. BC-Z does not have a terminate action, so we run these rollouts for a fixed maximum number of steps.
Key Questions and Results. We address the following questions in this work: (1) What is the success rate gap due to prompt embodiment (robot vs. human) across tasks? (§III-A1) (2) How do video-conditioned policies perform with unseen task videos? (Fig 5, §III-A2) (3) Is Vid2Robot’s overall success significantly better than BC-Z baseline? (§III-B) (4) Can learned motion representations handle out-of-distribution object interactions? (§III-C)
@@ III-A Task-based success
We compare our Vid2Robot model and baseline BC-Z with robot- and human-performed prompt videos in Table I. We train both Vid2Robot and BC-Z on the same data mixture containing robot-robot and human-robot paired data. Prompt videos cover a subset of the training tasks. However, the videos are unseen by the models. In this evaluation, we investigate each model’s ability to infer the task specification from the prompt video as well as the current observed state of the robot.
To test the model’s capabilities in different settings on real robots, we assess rollouts on the following nine tasks: ‘knock water bottle over’, ‘move rxbar chocolate near coke can’, ‘move green jalapeno chip bag near coke can’, ‘pick green rice chip bag’, ‘place coke can upright’, ‘pick coke can from bottom drawer and place on counter’, ‘open middle drawer’, ‘close middle drawer’, and ‘place apple into top drawer’.
We ask four evaluators to carry out two rollouts per task for a prompt video dataset and policy setting (a row in Table I). Overall, we have eight trials per task to evaluate a policy’s task success rate. We report an overall success rate per row over nine tasks with eight trials per task, that is, $9\!\times\!8\!=\!72$ trials. In total, we required $72\!\times\!4\!=\!288$ rollouts for Table I.
@@ III-A1 What is the gap in success rate due to embodiment difference in prompt videos?
When prompted with robot and human videos, we compare our model with BC-Z, which is a strong baseline for our comparisons. The overall success rate of our model Vid2Robot outperforms BC-Z for Human prompt videos by 20%, and is comparable for Robot prompt videos. Note that there is an order of magnitude more training samples for robot trajectories than human videos in our training mixture. Hence, there isn’t a significant gap in performance for robot prompt videos. Our model outperforms BC-Z in most tas for human prompt video, showing that Vid2Robot captures the task semantics from prompt videos better than the baseline. Our model outperforms in tasks like picking something from a drawer, placing it on the counter, and opening/closing drawers by a large margin. The most challenging task is placing upright and knocking over. We analyze the failure reasons in §V Fig 9.
III-A2 How well do video-conditioned policies perform when shown a task in an unseen video?
In addition to marking a rollout as a success, we recorded partial success annotations per rollout. In Fig 6, we observe that our model reaches to the correct object 78%, about 8% more than baseline. The policies sometimes fail to get the correct object and go towards a distractor instead. Next, grasping errors happen, particularly with small and deformable objects and collision-prone areas like drawer handles or counter’s edges. Here, our model (65%) outperforms BC-Z (45%) by a large margin of 20% — a successful grasp is often the most challenging part of a rollout and crucial for success. After grasping, most tasks require releasing at a correct location. Both models slightly drop in success rate due to incorrect release during the rollouts. While BC-Z runs for a fixed number of steps, our policy Vid2Robot predicts when to terminate. We observe that the rate of release and terminate is almost identical, about 57% for our model, which implies that after releasing at the correct location, Vid2Robot mostly terminates successfully.